OpenStudy (anonymous):

A car starts from rest and travels for 6.0 s with a uniform acceleration of +1.3 m/s2. The driver then applies the brakes, causing a uniform acceleration of -2.3 m/s2. The breaks are applied for 1.50 s. How fast is the car going at the end of the braking period? How far has the car gone from its start?

OpenStudy (anonymous):

For acceleration. v=at = 7.8 m/s displacement from origin = s1 = at^2/2 = 19.4 m For retardation v = u+at Here u=7.8 and a=-2.3 v=4.35 m/s displacement from s1 = s2= ut+at^2/2=9.1 m Displacemnt from origin = s1+s2=28.5m